[Ncep.nhc.nco_contacts] TAFB Shift Log -- Technical Edition -- Mon Sep 18, 2017

TAFB Shift Log nhc.tsbadmin at noaa.gov
Mon Sep 18 21:58:30 UTC 2017


TAFB Shift Log for Mon Sep 18 2017 17:58:27 GMT-0400 (EDT)

Shift: Swing Shift

Pacific Forecaster: C Landsea

Atlantic Forecaster: J Lewitsky

Surface Analyst: H Cobb

Technical Issues Summary:

GFE/NCP issues:

N-AWIPS issues:

Other issues: GOES-E imagery has had intermittent imagery with drop outs,  
horizontal black lines and blocks of missing imagery since yesterday.  
Called SDM and inquired if there was a reported issue nationally. They were  
unaware of this problem. HC tried to retrieve imagery using the "goes"  
command on Sheephead yesterday and script did not appear to terminate.  
Called Monica to have her investigate possible hung processes on Sheephead,  
as I could not open pgpallete there. Finally spoke with M DeMaria about  
issue during the afternoon and had him investigate to see if problem could  
be local.

Miscellaneous issues:
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170918/958c38a2/attachment-0001.html 


More information about the Ncep.nhc.nco_contacts mailing list